# Sampling and estimation on manifolds using the Langevin diffusion

### v2

@@ 1. Introduction
The Langevin algorithm for sampling from a target probability measure $\text{d}\mu_{\phi}\propto e^{-\phi(x)}\text{d}x$ on $\mathbb{R}^{q},q\geq 1$, is based on constructing a discrete time Markov process $\{X_{n}^{h},n\geq 1\}$ with time-step size $h=T/n>0$, $t_{n}=hn$, obtained by discretizing the ergodic Langevin diffusion
@@
In a statistical context, when sampling and estimation are of primary interest, quantitative error bounds, in terms of the time-step $h$ and integration time $T$, on two interconnected quantities related to distributional aspects of the approximations are relevant: bias of an estimator based on $X^{h}_{n}$ of the linear functional $\mu_{\phi}(\varphi):=\int\varphi\text{d}\mu_{\phi}$ over a class $\mathcal{H}$ of test functions $\varphi$; and an integral probability metric with respect to $\mathcal{H}$ between the stationary measure of $X_{n}^{h}$ and $\mu_{\phi}$. Such bounds are referred to as weak error bounds. In statistical applications (e.g., Bayesian computing), the commonly used estimator of $\mu_{\phi}(\varphi)$ is based on a single trajectory of $X^{h}_{n}$ until a large time $N$, known as the time-averaging estimator. Under suitable conditions on $\phi$, when the law of $X(t)$ converges to its invariant measure $\mu_{\phi}$ as $t\to\infty$, it is known that for the Euler method the optimal rate at which bias of the time-averaging estimator disappears is $O(h)$ (Talay, 1990; Milstein and Tretyakov, 2007; Mattingly et al., 2010; Milstein and Tretyakov, 2021); the first-order (in $h$) bound on the bias then results in a commensurate first-order bound on the integral probability metric.
@@
Modern statistical and machine learning applications have generated a growing need for extending the above program to a $q$-dimensional smooth Riemannian manifold $M$. To better understand the challenges, some context on the necessary ingredients is helpful. Broadly, nature of the applications engender two perspectives of, and hence coordinate systems on, $M$: (i) as a submanifold of $\mathbb{R}^{k},k\geq q$, with embedded geometry inherited from that of $\mathbb{R}^{k}$; (ii) as a $q$-dimensional differentiable manifold equipped with an intrinsic geometry. These two points of view lead to different formulations of the Langevin diffusion on $M$ with different invariant measures. The focus of this article is on the intrinsic perspective, and on investigating a geodesic-based sampling algorithm built on an intrinsic Langevin diffusion.
@@
A Riemannian metric $g$ prescribes a notion of curvature of the manifold $M$ at a point and determines a unique intrinsic Brownian motion on it (e.g., Elworthy, 1982; Hsu, 2002; Wang, 2013). The curvature of $M$ and behavior of the Brownian motion are intimately related: positive curvature makes curves come closer and thus prevents a Brownian motion from wandering away to infinity (non-explosion) while ensuring recurrence and fast convergence to ergodicity (e.g., Ichihara, 1982a, b); this is true even for its approximation based on a geodesic random walk (Mangoubi and Smith, 2018). Negative curvature, on the other hand, disperses curves and encourages Brownian motion to exit compact sets and drift away to infinity (explosion) while discouraging recurrence (e.g., Kendall, 1984). Ricci curvature at a point, which is, roughly, average of the sectional curvatures of two-dimensional subspaces of the tangent space at the point, captures the notion of curvature alluded to above. Compactness (unboundedness) and positive (non-positive) curvature of $M$ are closely related, although there are non-trivial exceptions depending on the Riemannian metric $g$ (for instance, manifolds with hyperbolic metric $g$ (Everitt and Maclachlan, 2000)). Thus, for negatively curved (typically) non-compact spaces a uniform lower bound on the Ricci curvatures is needed to ensure that the Brownian motion is well-behaved, while this is automatically true for positively curved (typically) compact manifolds. These issues, as one can now imagine, are subtler for a Brownian motion with drift.
@@
It would thus be reasonable to expect the intrinsic curvature of $M$, and regularity and behavior at infinity of its drift, to influence error of approximations of an intrinsically defined Langevin diffusion and its invariant measure. This is apparently the case in recent works (e.g., Wang et al., 2020; Gatmiry and Vempala, 2022; Cheng et al., 2022), where curvature-dependent constant terms show up in upper bounds on distances between the stationary measure $\mu^{h}_{\phi}$ of a discretized Markov processes $X^{h}_{n}$ and $\mu_{\phi}$ on $M$ defined with respect to the Riemannian volume measure $\mathrm{dvol}_{g}$. Importantly, however, the curvature-terms do not affect the claimed error order with respect to step size $h$.
@@
A differentiable manifold $M$ is locally Euclidean, covered by a family of charts (open subsets of $\mathbb{R}^{q}$) and smooth mappings between them. In a small region of $M$ local analysis of the weak error within a chart, where $X$ solves the appropriate stochastic differential equation (SDE), proceeds as in $\mathbb{R}^{q}$ using the local coordinate representation of the metric $g$ in the chart, which accounts for the local curvature. The challenge lies in adapting the analysis when one transitions to another chart, where the local representation of the metric $g$, and hence that of $X$, changes. Thus, on flat manifolds $M$ with zero curvature, error analysis proceeds unabated as in $(\mathbb{R}^{q},g)$. Indeed, Mattingly et al. (2010) demonstrate precisely this with weak error analysis when $M=\mathbb{T}^{q}\cong\mathbb{R}^{q}/\mathbb{Z}^{q}$, the $q$-dimensional compact torus, equipped with the flat (quotient) metric from $\mathbb{R}^{q}$. The main technical ingredient in their analysis was the link between the solution of a Poisson equation (given by a partial differential equation (PDE)) on $\mathbb{T}^{q}$ and the infinitesimal generator of the Langevin diffusion $X$. Aided by the fact that a weak analysis merely requires bounding and matching of a handful of moments, assured by compactness of $\mathbb{T}^{q}$, an optimal first-order $O(h)$ weak error bound, matching the situation in $\mathbb{R}^{q}$, was established for general discretization schemes, including the Euler method.
@@ 1.1. Contributions and related work
Specifically, on a connected compact Riemannian manifold $(M,g)$ without boundary we consider an intrinsically defined Langevin diffusion
@@
Two estimators of the linear functional $\mu_{\phi}(\varphi)$ for a smooth $\varphi$ are proposed: an ensemble-averaging estimator defined using several independent trajectories of $X^{h}_{n}$, and a time-averaging estimator based on a single trajectory. The former is of great practical merit. First, it enables decoupling and management of three sources of error: (i) length of the integration time $T$ that controls proximity of $X$ to its ergodic limit; (ii) discretization error controlled by the step size $h$; (iii) the number of independent trajectories of $X^{h}_{n}$ that controls the Monte Carlo error (sample variance). Second, the estimator is particularly useful in a parallel computing environment (GPUs or array of CPUs) where independent trajectories of $X^{h}_{n}$ can be generated in parallel and combined in the end. Analysis of the time-averaging estimator leads to first-order bound on a distance between the empirical measure $\mu^{h}_{\phi}$ of $X^{h}_{n}$ and $\mu_{\phi}$, with respect to a smooth class $\mathcal{H}$ of test functions.
@@
In existing works on SDE-based algorithms for sampling on manifolds (Gatmiry and Vempala, 2022; Li, 1992; Cheng et al., 2022) weak convergence of the algorithm is established by first proving mean-square (strong) convergence, since weak convergence follows from mean-square; a consequence of this approach is a non-optimal weak convergence order when compared to the situation in $\mathbb{R}^{q}$. A notable exception to this is Grorud and Talay (1996), which we discuss in Section 3.1.
@@
Thanks to the PDE-based technique for proving weak convergence (Milstein, 1985; Talay, 1986, 1990; Mattingly et al., 2010) (see also Milstein and Tretyakov (2021)), our proofs of weak convergence for Riemannian Langevin algorithms have a unified nature and are easily adaptable to a variety of schemes, for example, involving retraction maps (see Section 3.2). The consequence of this proof technique is that we do not monitor dependence of the coefficients (at $h$) in upper error bounds on parameters of the problem (such as dimension, curvature, spectral gap, etc). However, using the Talay-Tuboro expansion (Talay and Tubaro, 1990) (see Section 4.4) one can estimate the leading error term of a weak scheme (i.e., the posteriori error), which is valuable in practice.
@@
As mentioned earlier, there are some existing works that provide curvature-dependent bounds for distances between $\mu^{h}_{\phi}$ and $\mu_{\phi}$ under different intrinsic discretizations $X^{h}_{n}$ and settings: Grorud and Talay (1996) for chart-dependent Euler scheme (see further discussion at the end of Section 3.1); Cheng et al. (2022) for the geodesic Euler scheme considered here claim to develop $O(h^{1/2})$ bound in expected Riemannian distance between $X(t)$ and $X_{n}^{h}$, $t=nh$; Gatmiry and Vempala (2022) provide an $O(h)$ bound on the Kullback-Leibler distance for Hessian manifolds, but assume possibility of exact sampling of a Brownian increment on $M$, which currently is available only for spheres or products of sphere (Li, 1992). Further, the above works do not consider and analyse an estimator of $\mu_{\phi}(\varphi)$ while we obtain an $O(h)$ bound on the bias of the estimators.
@@
Although not directly related to our work, we note that there are quite a few papers on extrinsic discretization methods for an embedded Langevin diffusion $X$ defined by viewing $M$ as a submanifold of $\mathbb{R}^{k},k\geq q$ (e.g., Leliévre et al., 2012; Sharma and Zhang, 2021; Laurent and Vilmart, 2022; Armstrong and King, 2022). A critical limitation of extrinsic methods is that while $X$ is constrained to move on $M$, the discretized Markov process $X^{h}_{n}$ need not, and repeated, computationally expensive, projections on to $M$ are required to ensure that $X^{h}_{n}$ does not leave $M$; an exception is when $M$ is a Lie group, and the group action enables transport along the manifold (e.g., Malham and Wiese, 2008; Davidchack et al., 2009; Mentink et al., 2010; Davidchack et al., 2015). Use of projections results in an additional source of error over the original discretization of $X$, avoided entirely by intrinsic methods. Intrinsic methods may require additional computation of geodesics when they are unknown in closed-form, but this can be addressed by using computionally cheaper, higher-order retractions which do not affect the theoretical weak error bounds (see Section 3.2).
@@ 2. Preliminaries
In Section 2.1 we briefly review concepts and fix notation from differential geometry needed for analysis on a Riemannian manifold $M$; for more details we refer to some well-known sources (e.g., Jost, 2008; Cheeger and Ebin, 1975; Gallot et al., 1990; Bishop and Crittenden, 2011). In Section  2.2 we introduce the Langevin diffusion on $M$ with a given invariant measure, the target measure for sampling.
@@ 2.1. Differential geometric concepts
Tangent spaces $T_{x}M$ and tangent bundle $TM$ have dual spaces, the cotangent space $T_{x}^{*}M$ and cotangent bundle $T^{*}M$ consisting of linear functionals $T_{x}M\to\mathbb{R}$. The coordinate basis $\{\partial_{x^{1}},\ldots,\partial_{x^{d}}\}$ induces a similar basis $\{\text{d}x^{1},\ldots,\text{d}x^{q}\}$ on $T^{*}_{x}M$ defined via the evaluation $\text{d}x^{j}(\partial_{x^{i}})=\delta^{j}_{i}$, where $\delta^{j}_{i}$ is the delta function, such that $\nu=\nu_{i}\text{d}x^{i}$ for every $\nu\in T^{*}_{x}M$.
A differentiable manifold $M$ becomes a Riemannian manifold $(M,g)$ when equipped with a Riemannian metric $g:T_{x}M\times T_{x}M\to\mathbb{R}$, a family $\{g(x),x\in M\}$ of symmetric positive definite bilinear maps (inner products) which varies smoothly on $M$. In a chart around $x$, the metric $g(x)$ can be expressed as a symmetric positive definite matrix whose $ij$th entry is $g(\partial_{x^{i}},\partial_{x^{j}})$.
Denote by $D$ the unique torsion-free Levi-Civita connection, or covariant derivative, compatible with $g$. The connection enables definition of a derivative $D_{X}Y$ of a vector field $X$ along a vector field $Y$. For a smooth function $f:M\to\mathbb{R}$, denote by $X(f)$ the derivative of $f$ in the direction of $X$ based on the interpretation of tangent vectors as derivations. The covariant derivative is linear in $X$ and follows the product rule in $Y$ so that $D_{X}(fY)=X(f)Y+fD_{X}Y$. More conveniently, $D_{X}Y$ may be written in local coordinates at $x$ using the relation $D_{\partial_{x^{i}}}\partial_{x^{j}}=\Gamma^{k}_{ij}\partial_{x^{k}}$, where $\Gamma^{k}_{ij}$ are the Christoffel symbols.
@@
It is necessary to introduce a handful of tensors in preparation for the analysis of the algorithm. First, the exterior derivative d is a differential operator that maps $C^{1}$ functions to the cotangent space $T^{*}M$, and in local coordinates assumes the form $\text{d}f(x)=\frac{\partial f(x)}{\partial x^{i}}\text{d}x^{i}$. The quantity $\text{d}f$ is known as a 1-form. More generally, an $r$-form, for $r\in\{1,\ldots,q\}$, can be considered on the $\wedge^{r}T_{x}^{*}M$ with anti-symmetric wedge product $\wedge$ spanned by $\text{d}x^{1}\wedge\cdots\wedge\text{d}x^{r}$. For brevity, the wedge symbol is omitted when discussing differential forms. Then, the exterior derivative d maps $r$-forms to $(r+1)$-forms and $\text{d}^{2}=0$. For the purposes of integration, we introduce the volume form, a $q$-form, $\mathrm{dvol}_{g}=\sqrt{|g|}\text{d}x^{1}\ldots\text{d}x^{q}$, where $|g|$ denotes the determinant of the Riemannian metric $g$ in local coordinates around $x$.
@@
defines the Riemannian gradient operator $\nabla$ acting on smooth functions $f:M\to\mathbb{R}$ as the dual of $\text{d}f$. In local coordinates, the gradient is written as $\nabla f=g^{ij}\frac{\partial f}{\partial x^{i}}\frac{\partial}{\partial x^{j}}$; it is often convenient to omit dependence on $x$ and use $\text{d}f(v),v\in T_{x}M$.
@@
Then, the Laplace–Beltrami operator, as an analogue of the Laplacian in $\mathbb{R}^{q}$, is defined as the trace (with respect to the metric tensor) of the Hessian
$\Delta f=\mathrm{Tr}\mathrm{Hess}^{f}=g^{ij}\mathrm{Hess}_{ij}^{f}.$
@@
Define the operator norm for a $(0,k)$-tensor $\mathfrak{T}$ on $M$
@@
For vector fields $X,Y,Z$, the Riemannian curvature tensor $R:TM\times TM\times TM$of $M$ is defined as
@@
where is the commutativity bracket defined via its application to a smooth function $f:M\to\mathbb{R}:[X,Y]f=X(Y(f))-Y(X(f))$. Roughly, it quantifies the difference between the vectors obtained by transporting a vector $Z(x)$ first along $X(x)$ then $Y(x)$, in contrast to moving first along $Y(x)$ and then $X(x)$. The sectional curvature at $x$ is the quadratic form (Hsu, 2002, pp. 89)
@@
$\mathrm{Ric}(X,Y):=\sum_{i=1}^{q}g(R(X,X_{i})X_{i},X).$

### v3

@@ 1. Introduction
The Langevin algorithm for sampling from a target probability measure $\text{d}\mu_{\phi}\propto e^{-\phi(x)}\text{d}x$ on $\mathbb{R}^{q},q\geq 1$, is based on constructing a discrete time Markov process $\{X_{n}^{h},n\geq 1\}$ with time-step size $h=T/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N}>0$, $t_{n}=hn$, obtained by discretizing the ergodic Langevin diffusion
@@
In a statistical context, when sampling and estimation are of primary interest, quantitative error bounds, in terms of the time-step $h$ and integration time $T$, on two interconnected quantities related to distributional aspects of the approximations are relevant: bias of an estimator based on $X^{h}_{n}$ of the linear functional $\mu_{\phi}(\varphi):=\int\varphi\text{d}\mu_{\phi}$ over a class $\mathcal{H}$ of test functions $\varphi$; and an integral probability metric with respect to $\mathcal{H}$ between the stationary measure of $X_{n}^{h}$ and $\mu_{\phi}$. Such bounds are referred to as weak error bounds. In statistical applications (e.g., Bayesian computing), the commonly used estimator of $\mu_{\phi}(\varphi)$ is based on a single trajectory of $X^{h}_{n}$ until a large time $T=Nh$, known as the time-averaging estimator. Under suitable conditions on $\phi$, when the law of $X(t)$ converges to its invariant measure $\mu_{\phi}$ as $t\to\infty$, it is known that for the Euler method the optimal rate at which bias of the time-averaging estimator disappears is $O(h)$ (Talay, 1990; Milstein and Tretyakov, 2007; Mattingly et al., 2010; Milstein and Tretyakov, 2021); the first-order (in $h$) bound on the bias then results in a commensurate first-order bound on the integral probability metric.
@@
Modern statistical and machine learning applications have generated a growing need for extending the above program to a $q$-dimensional smooth Riemannian manifold $M$. To better understand the challenges, some context on the necessary ingredients is helpful. Broadly, the type of application engenders two perspectives of, and hence coordinate systems on, $M$: (i) as a submanifold of $\mathbb{R}^{k},k\geq q$, with embedded geometry inherited from that of $\mathbb{R}^{k}$; (ii) as a $q$-dimensional differentiable manifold equipped with an intrinsic geometry. These two points of view lead to different formulations of the Langevin diffusion on $M$ with different invariant measures. The focus of this article is on the intrinsic perspective, and on investigating a geodesic-based sampling algorithm built on an intrinsic Langevin diffusion.
@@
A Riemannian metric $g$ prescribes a notion of curvature of the manifold $M$ at a point and determines a unique intrinsic Brownian motion on it (e.g., Elworthy, 1982; Hsu, 2002; Wang, 2013). The curvature of $M$ and behavior of the Brownian motion are intimately related: positive curvature makes curves come closer and thus prevents a Brownian motion from wandering away to infinity (non-explosion) while ensuring recurrence and fast convergence to ergodicity (e.g., Ichihara, 1982a, b); this is true even for its approximation based on a geodesic random walk (Mangoubi and Smith, 2018). Negative curvature, on the other hand, disperses curves and encourages Brownian motion to exit compact sets and drift away to infinity (explosion) while discouraging recurrence (e.g., Kendall, 1984). Ricci curvature at a point, which is, roughly, the average of the sectional curvatures of two-dimensional subspaces of the tangent space at the point, captures the notion of curvature alluded to above. Compactness (unboundedness) and positive (non-positive) curvature of $M$ are closely related, although there are non-trivial exceptions depending on the Riemannian metric $g$ (for instance, manifolds with hyperbolic metric $g$ (Everitt and Maclachlan, 2000)). Thus, for negatively curved (typically) non-compact spaces , a uniform lower bound on the Ricci curvature is needed to ensure that the Brownian motion is well-behaved, while this is automatically true for positively curved (typically) compact manifolds. These issues, as one can now imagine, are subtler for a Brownian motion with drift.
@@
It would thus be reasonable to expect the intrinsic curvature of $M$, and regularity and behavior at infinity of its drift, to influence error of approximations of an intrinsically defined Langevin diffusion and its invariant measure. This is apparently the case in recent works (e.g., Gatmiry and Vempala, 2022; Cheng et al., 2022), where curvature-dependent constant terms show up in upper bounds on distances between the stationary measure $\mu^{h}_{\phi}$ of a discretized Markov processes $X^{h}_{n}$ and $\mu_{\phi}$ on $M$ defined with respect to the Riemannian volume measure $\mathrm{dvol}_{g}$. Importantly, however, the curvature-terms do not affect the claimed error order with respect to step size $h$.
@@
A differentiable manifold $M$ is locally Euclidean, covered by a family of charts (open subsets of $\mathbb{R}^{q}$) and smooth mappings between them. In a small region of $M$ , local analysis of the weak error within a chart, where $X$ solves the appropriate stochastic differential equation (SDE), proceeds as in $\mathbb{R}^{q}$ using the local coordinate representation of the metric $g$ in the chart, which accounts for the local curvature. The challenge lies in adapting the analysis when one transitions to another chart, where the local representation of the metric $g$, and hence that of $X$, changes. Thus, on flat manifolds $M$ with zero curvature, error analysis proceeds unabated as in $(\mathbb{R}^{q},g)$. Indeed, Mattingly et al. (2010) demonstrate precisely this with weak error analysis when $M=\mathbb{T}^{q}\cong\mathbb{R}^{q}/\mathbb{Z}^{q}$, the $q$-dimensional compact torus, equipped with the flat (quotient) metric from $\mathbb{R}^{q}$. The main technical ingredient in their analysis was the link between the solution of a Poisson equation (given by a partial differential equation (PDE)) on $\mathbb{T}^{q}$ and the infinitesimal generator of the Langevin diffusion $X$. Aided by the fact that a weak error analysis merely requires bounding and matching of a handful of moments, assured by compactness of $\mathbb{T}^{q}$, an optimal first-order $O(h)$ weak error bound, matching the situation in $\mathbb{R}^{q}$, was established for general discretization schemes, including the Euler method.
@@ 1.1. Contributions and related work
Specifically, on a connected compact Riemannian manifold $(M,g)$ without boundary , we consider an intrinsically defined Langevin diffusion
@@
Two estimators of the linear functional $\mu_{\phi}(\varphi)$ for a smooth $\varphi$ are proposed: an ensemble-averaging estimator defined using multiple independent trajectories of $X^{h}_{n}$, and a time-averaging estimator based on a single trajectory. The former is of great practical merit. First, it enables decoupling and management of three sources of error: (i) length of the integration time $T$ that controls proximity of $X$ to its ergodic limit; (ii) discretization error controlled by the step size $h$; (iii) the number of independent trajectories of $X^{h}_{n}$ that controls the Monte Carlo error (sample variance). Second, the estimator is particularly useful in a parallel computing environment (GPUs or array of CPUs) where independent trajectories of $X^{h}_{n}$ can be generated in parallel and combined in the end. Analysis of the time-averaging estimator leads to first-order bound on a distance between the empirical measure $\mu^{h}_{\phi}$ of $X^{h}_{n}$ and $\mu_{\phi}$, with respect to a smooth class $\mathcal{H}$ of test functions.
@@
In existing works on SDE-based algorithms for sampling on manifolds (Gatmiry and Vempala, 2022; Cheng et al., 2022) weak convergence of the algorithm is established by first proving mean-square (strong) convergence, since weak convergence follows from mean-square; a consequence of this approach is usually a non-optimal weak convergence order when compared to the situation in $\mathbb{R}^{q}$. A notable exception to this is Grorud and Talay (1996), which we discuss in Section 3.1. Mean-square convergence of numerical methods for intrinsically defined SDEs on a Riemannian manifold is studied in Bharath et al. (2025).
@@
Thanks to the PDE-based technique for proving weak convergence (Milstein, 1985; Talay, 1986, 1990; Mattingly et al., 2010) (see also Milstein and Tretyakov (2021)), our proofs of weak convergence for Riemannian Langevin algorithms have a unified nature and are easily adaptable to a variety of schemes, for example, involving retraction maps (see Section 3.2). The consequence of this proof technique is that we do not monitor dependence of the coefficients (at $h$) in upper error bounds on parameters of the problem (such as dimension, curvature, spectral gap, etc). However, using the Talay- Tubaro expansion (Talay and Tubaro, 1990) (see Section 4.4) , one can estimate the leading error term of a weak scheme (i.e., the posteriori error), which is valuable in practice.
@@
As mentioned earlier, there are some existing works that provide curvature-dependent bounds for distances between $\mu^{h}_{\phi}$ and $\mu_{\phi}$ under different intrinsic discretizations $X^{h}_{n}$ and settings: Grorud and Talay (1996) for chart-dependent Euler scheme (see further discussion at the end of Section 3.1); Cheng et al. (2022) for the geodesic Euler scheme considered here claim to develop $O(h^{1/2})$ bound in expected squared Riemannian distance between $X({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t_{n}})$ and $X_{n}^{h}$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t_{n}}=nh$; Gatmiry and Vempala (2022) provide an $O(h)$ bound on the Kullback-Leibler distance for Hessian manifolds, but assume possibility of exact sampling of a Brownian increment on $M$, which currently is available for spheres (Mijatović et al., 2020). Further, the above works do not consider and analyse an estimator of $\mu_{\phi}(\varphi)$ while we obtain an $O(h)$ bound on the bias of the estimators.
@@
Although not directly related to our work, we note that there are quite a few papers on extrinsic discretization methods for an embedded Langevin diffusion $X$ defined by viewing $M$ as a submanifold of $\mathbb{R}^{k},k\geq q$ (e.g., Leliévre et al., 2012; Sharma and Zhang, 2021; Laurent and Vilmart, 2022; Armstrong and King, 2022). A critical limitation of extrinsic methods is that while $X$ is constrained to move on $M$, the discretized Markov process $X^{h}_{n}$ need not, and repeated, computationally expensive, projections on to $M$ are required to ensure that $X^{h}_{n}$ does not leave $M$; an exception is when $M$ is a Lie group, and the group action enables transport along the manifold (e.g., Malham and Wiese, 2008; Davidchack et al., 2009; Mentink et al., 2010; Davidchack et al., 2015). Use of projections results in an additional source of error over the original discretization of $X$, avoided entirely by intrinsic methods. Intrinsic methods may require additional computation of geodesics when they are unknown in closed-form, but this can be addressed by using computationally cheaper, higher-order retractions which do not affect the theoretical weak error bounds (see Section 3.2). We note that as with the approximation of the projection map in the extrinsic setting, an efficient computation of retraction in our intrinsic setting does not depend on the potential $\phi$ and noise, and depends only on the manifold.
Projection-based schemes are inappropriate when the target probability density to sample from on a manifold $M$ is defined with respect to a volume form coming from a specific Riemannian metric $g$ that is not the restriction of the standard metric in $\mathbb{R}^{k},k\geq q$. Moreover, when $M$ is an open submanifold of $\mathbb{R}^{k}$, in the absence of a level-set representation of $M$, projection-based schemes are inappropriate. A practically relevant example that highlights the above two issues concerns the manifold $\mathcal{P}_{m}$ of $m\times m$ real symmetric positive definite matrices of dimension $q=m(m+1)/2$ (see Section 6.2). Firstly, it is an open cone within the same dimensional vector space of symmetric matrices so that $k=m(m+1)/2$; an orthogonal projection would map a symmetric matrix with a single negative eigenvalue to a positive definite matrix with smallest eigenvalue very close to zero, close to the boundary of $\mathcal{P}_{m}$, which is infinitely far away with respect to the intrinsic geometry of $\mathcal{P}_{m}$, and the projection-based sampling scheme will fail to sample from the target measure with support in $\mathcal{P}_{m}$. Secondly, there are numerous choices of intrinsic Riemannian metrics motivated by practical applications (e.g. Pennec et al., 2006, Chapter 3) which lead to specific target probability densities that cannot be sampled with projection-based schemes that views $M$ as a submanifold of $\mathbb{R}^{k}$ with the induced metric.
Additionally, projection-based schemes are typically less stable (requiring the use of appreciably smaller time steps) compared to geometric integrators, which by construction automatically lie on the manifold; see, for example, Mentink et al. (2010) when $M=\mathbb{S}^{2}$. For $M$ that is a Lie group or a homogeneous space of one, Lie group integrators preserving the geometry represent an efficient choice. However, for general manifolds $M$ with specific metrics $g$ the intrinsic methods considered in this paper will typically have an advantage when compared with projection-based schemes.
@@ 2. Preliminaries
In Section 2.1 , we briefly review concepts and fix notation from differential geometry needed for analysis on a Riemannian manifold $M$; for more details we refer to some well-known sources (e.g., Jost, 2008; Cheeger and Ebin, 1975; Gallot et al., 1990; Bishop and Crittenden, 2011). In Section 2.2 , we introduce the Langevin diffusion on $M$ with a given invariant measure, the target measure for sampling.
@@ 2.1. Differential geometric concepts
Tangent spaces $T_{x}M$ and the tangent bundle $TM$ have dual spaces, the cotangent space $T_{x}^{*}M$ and cotangent bundle $T^{*}M$ consisting of linear functionals ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\nu}:T_{x}M\to\mathbb{R}$. The coordinate basis $\{\partial_{x^{1}},\ldots,\partial_{x^{d}}\}$ induces a similar basis $\{\text{d}x^{1},\ldots,\text{d}x^{q}\}$ on $T^{*}_{x}M$ defined via the evaluation $\text{d}x^{j}(\partial_{x^{i}})=\delta^{j}_{i}$, where $\delta^{j}_{i}$ is the delta function, such that $\nu=\nu_{i}\text{d}x^{i}$ for every $\nu\in T^{*}_{x}M$.
A differentiable manifold $M$ becomes a Riemannian manifold $(M,g)$ when equipped with a Riemannian metric $g:T_{x}M\times T_{x}M\to\mathbb{R}$, a family $\{g(x),x\in M\}$ of smoothly varying symmetric positive definite bilinear forms (inner products). In a chart around $x$, the metric $g(x)$ can be expressed as a symmetric positive definite matrix whose $(i,j)$th entry is $g(\partial_{x^{i}},\partial_{x^{j}})$, which we write in index notation as $g_{ij}$.
Denote by $D$ the unique torsion-free Levi-Civita connection, or covariant derivative, compatible with $g$. The connection enables definition of a derivative $D_{X}Y$ of a vector field ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y}$ along a vector field ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X}$. For a smooth function $f:M\to\mathbb{R}$, denote by $X(f)$ the derivative of $f$ in the direction of $X$ based on the interpretation of tangent vectors as derivations. The covariant derivative is linear in $X$ and follows the product rule in $Y$ so that $D_{X}(fY)=X(f)Y+fD_{X}Y$. More conveniently, $D_{X}Y$ may be written in local coordinates at $x$ using the relation $D_{\partial_{x^{i}}}\partial_{x^{j}}=\Gamma^{k}_{ij}\partial_{x^{k}}$, where $\Gamma^{k}_{ij}$ are the Christoffel symbols.
@@
It is necessary to introduce a handful of tensors in preparation for the analysis of the algorithm. First, the exterior derivative d is a differential operator that maps $C^{1}$ functions to the cotangent bundle $T^{*}M$, and in local coordinates assumes the form $\text{d}f(x)=\frac{\partial f(x)}{\partial x^{i}}\text{d}x^{i}$. The quantity $\text{d}f$ is known as a 1-form. More generally, an $r$-form, for $r\in\{1,\ldots,q\}$, can be considered on $\wedge^{r}T_{x}^{*}M$ with anti-symmetric wedge product $\wedge$ spanned by $\text{d}x^{1}\wedge\cdots\wedge\text{d}x^{r}$. For brevity, the wedge symbol is omitted when discussing differential forms. Then, the exterior derivative d maps $r$-forms to $(r+1)$-forms and $\text{d}^{2}=0$. For the purposes of integration, we introduce the volume form, a $q$-form, $\mathrm{dvol}_{g}=\sqrt{|g|}\text{d}x^{1}\ldots\text{d}x^{q}$, where $|g|$ denotes the determinant of the Riemannian metric $g$ in local coordinates around $x$.
@@
defines the Riemannian gradient operator $\nabla$ acting on smooth functions $f:M\to\mathbb{R}$ as the dual of $\text{d}f$. In local coordinates, the gradient is written as $\nabla f=g^{ij}\frac{\partial f}{\partial x^{i}}\frac{\partial}{\partial x^{j}}$ where $g^{ij}$ are the elements of the inverse metric; it is often convenient to omit dependence on $x$ and use $\text{d}f(v),v\in T_{x}M$.
@@
Then, the Laplace-Beltrami operator, as an analogue of the Laplacian in $\mathbb{R}^{q}$, is defined as the trace (with respect to the metric tensor) of the Hessian
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta_{M}}f=\mathrm{Tr}\mathrm{Hess}^{f}=g^{ij}\mathrm{Hess}_{ij}^{f}.$
@@
Define the operator norm for a $(0,k)$-tensor $\mathfrak{T}$ on $M$ (e.g. Le et al., 2024)
@@
For vector fields $X,Y,Z$, the Riemannian curvature tensor at a point $x\in M$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R:T_{x}M\times T_{x}M\times T_{x}M\rightarrow T_{x}M}$ of $M$ is defined as
@@
where the commutativity bracket (or Lie bracket) is defined via its application to a smooth function $f:M\to\mathbb{R}$ by $[X,Y]f=X(Y(f))-Y(X(f))$. Roughly, it quantifies the difference between the vectors obtained by transporting a vector $Z(x)$ first along $X(x)$ then $Y(x)$, in contrast to moving first along $Y(x)$ and then $X(x)$. The sectional curvature at $x$ is the quadratic form (Hsu, 2002, pp. 89)
@@
$\mathrm{Ric}(X,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X}):=\sum_{i=1}^{q}g(R(X,X_{i})X_{i},X).$